A 3.5 kg rifle fires an 11g bullet with a velocity of 575 m/s. a.) What is the recoil velocity of the rifle. b.) If the shooter holds the rifle firmly against his shoulder,what will be the recoil velocity?


Assume the mass of shooter is 45kg

1
Expert's answer
2022-01-18T10:24:33-0500

a) Find the recoil velocity:


"0=mv+Mu→u=v\dfrac mM=1.81\text{ m/s}."


b) In this case:


"0=mv+(M+M_s)u→v\dfrac{m}{M+M_s}=0.13\text{ m/s}."